# Beyond Pipelines: A Survey of the Paradigm Shift toward Model-Native Agentic AI

### v1

@@ 1.1. Paradigms
  • Memory has long been addressed with external modules. Short-term memory was typically maintained through conversation summary (Chase and Gola, 2023), where long interaction histories were summarized and reinserted to fit within the context window. Long-term memory usually relied on retrieval-based methods, most prominently through Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), which stored past interactions in a vector database and retrieved them on demand.
@@
  • Memory has also begun to shift from external modules to model-native mechanisms. For short-term memory, Qwen-2.5-1M (Yang et al., 2025e) leveraged synthetic long-sequence data to extend the native context windows, addressing the challenge of “remembering” long-horizon information. Beyond remembering, recent works attempt to effectively “using” the information, e.g., MemAct (Zhang et al., 2025c) reframed context management as a tool the agent learns to call, proactively deciding when and how to store or retrieve information based on dynamic state and environmental feedback. For long-term memory, representative works like MemoryLLM (Wang et al., 2024a) pioneered parameterizing memory directly, where a set of latent memory tokens is continuously updated as part of the model’s forward pass, resulting in automatically updated internal knowledge.
@@ 1.2. Applications
The paradigm shifted with OpenAI’s introduction of the first model-native Deep Research agent (OpenAI, [n. d.]), which was fine-tuned based on its agentic foundation model o3. This approach, where the model’s internal policy learns to strategize the entire research process, significantly enhanced long-horizon consistency and the depth of information discovery, establishing Deep Research as an advanced assistant capable of tackling rigorous professional tasks. More recently, Tongyi Lab’s WebAgent series have advanced the open-source implementation of model-native Deep Research agents. This includes WebSailor (Li et al., 2025j), which addressed the critical challenge of synthesizing high-quality trajectory data for uncertain tasks, and culminated in the Tongyi DeepResearch model (Lab, 2025), which is capable of executing complex, multi-step research tasks in dynamic web environments.
@@ 1.3. Algorithms
To address these issues, a new set of outcome-driven RL algorithms has emerged to address the practical challenges of training stability and efficiency in long-horizon tasks. To counter the inefficiency of PPO’s large critic network, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) was proposed. It introduces a lightweight evaluation paradigm that computes advantages based on the relative rewards within a group of sampled responses, circumventing the need for an absolute value critic and improving training stability. Further refining this, Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO) (Yu et al., 2025) improves performance in multi-turn interactions by decoupling the clipping mechanism for positive and negative advantages and employing dynamic sampling strategies, making it particularly effective for training long-horizon agents.
@@ 2.1. Necessity: The Shortage of Procedural Data
Methodological Singularity Shift in Research Focus
Application Expansion Capability Enhancement Theoretical Innovation
Physics Newtonian Mechanics + Calculus Fluid dynamics, Celestial mechanics Many-Body, Turbulence problem Electromagnetic Theory, Quantum
AI LLM + RL Scientific Discovery, Social Systems Continual Learning, Safety & Alignment -
Table 1. Physics vs. AI: methodological singularities and subcequent shifts in research focus.
@@
where $\mathcal{M}_{\text{base}}$ provides general world knowledge and reasoning priors, $\mathcal{A}_{\text{learn}}$ (e.g., RL or preference optimization) adapts and refines these capabilities through interaction and optimization, and $\mathcal{E}_{\text{task}}$ defines the environment, tool set, and reward signals that contextualize learning. Recently, a growing body of work demonstrates how research across NLP, vision, robotics, and scientific domains is increasingly converging on this formula.
This is strikingly analogous to the unification of Physics after Newton. Before Newton, physical subfields were fragmented: Celestial mechanics was governed by Kepler’s laws describing planetary motion, Terrestrial mechanics was pioneered by Galileo’s studies of falling bodies, and fields like Optics and Fluid dynamics had their own disparate empirical and geometric principles. Newton’s “Three Laws of Motion”, “Law of Universal Gravitation”, together with the powerful mathematical tool of Calculus, provided a universal framework that unified the experimental research of the heavens and the earth under a single set of principles, achieving an unprecedented unification for Physics.
This precisely resembles the role that “LLM + RL” plays today for AI. LLM provides a unified model of world knowledge and foundational reasoning, much like Newton’s laws. Meanwhile, RL offers a dynamic, goal-oriented optimization framework, analogous to the problem-solving engine of Calculus. Following the development of Physics after Newton, as summarized in Table 1, we discuss the shift in AI’s research focus after the “LLM + RL” methodological singularity. The new focus may range from the expansion of applications into practical and complex domains like healthcare, scientific discovery and social system simulation, to the enhancement of advanced agentic capabilities such as continual learning, safety and alignment, and ultimately toward possibly new theories on general intelligence.
@@ 3.3.1. Supervised Learning
  • Tree search.  These typically combine Monte Carlo estimation or process reward models to enable fine-grained control over step-level quality (Li et al., 2025h, b). For instance, Marco-o1 (Zhao et al., 2024) employs Monte Carlo Tree Search (MCTS), using model confidence to guide exploration and synthesize high-quality reasoning data. WebSynthesis (Gao et al., 2025b) introduces a world model in Web-UI agent scenarios, combined with MCTS to generate web interaction trajectories at scale, supported by a two-stage curriculum learning strategy to enhance policy capability. ReST-MCTS* (Zhang et al., 2024i) integrates a process reward model with MCTS to collect higher-quality trajectories by assessing the value of each reasoning step. Regarding multimodal reasoning, AStar (Wu et al., 2025b) defines an executable visual action space and leverages MCTS to obtain high-quality visual reasoning data.
@@ 3.3.2. Reinforcement Learning
    To address scalability limitations of manual labeling, many studies employ Monte Carlo (MC) estimation or LLM-as-Judge (Li et al., 2025i) methods, sampling and comparing multiple reasoning paths to assess step quality. For example, Math-Shepherd (Wang et al., 2024b) assigns correctness scores to each mathimatical inference step via MC estimation, eliminating dependency on human annotation. OmegaPRM (Luo et al., 2024) enhances this approach by combining MC estimation with Monte Carlo Tree Search (MCTS), using binary search to locate the first erroneous step in reasoning chains, thereby improving data collection efficiency. Qwen-2.5-Math-PRM (Zhang et al., 2025f) notes that both MC estimation and LLM-as-Judge introduce noise in automatic labeling, while human annotation remains cost-prohibitive, thus proposing a hybrid approach to reduce noise and improve accuracy.
@@
Outcome reward, also referred to as outcome supervision, provides reward signals based solely on the correctness of the final answer. It offers a sparse yet unambiguous reward signal, grounded in the core hypothesis that a powerful foundation model can, through feedback on end results, autonomously explore and internalize effective reasoning policy. Exemplified by the RLVF approach adopted in DeepSeek R1 (DeepSeek-AI, 2025), outcome reward completely bypasses the complex evaluation of intermediate steps and only requires verifying the correctness of the final answer (Team, 2025b; He et al., 2025a). In tasks with easily verifiable answers, such as multiple-choice questions and mathematical problems, RL is typically applied using verifiable outcome rewards to enhance the model’s reasoning capability (Trung et al., 2024; Zhang et al., 2024g; Luo et al., 2025c). Similarly, in the domain of code generation, methods like o1-Coder (Zhang et al., 2024d) and RLEF (Gehring et al., 2025) validate answer correctness based on program execution feedback.
@@ 3.4. Summary and Discussion
  • From supervised to self-supervised/unsupervised internalization: To reduce reliance on external reward signals, future research will explore self-supervised or unsupervised learning mechanisms. For example, by designing intrinsic reward signals such as exploration rewards based on information-theoretic principles like information gain or entropy (Zhang et al., 2025d), models could be motivated to autonomously discover novel and complex reasoning strategies, enabling continuous learning without explicit external objectives.
  • From single-task to general and transferable planning: Current models are often trained for planning within specific domains (e.g., mathematics or code). A significant future challenge lies in achieving cross-domain generalization and transfer of planning abilities (Su et al., 2025; Liu et al., 2025c). This requires models not only to learn problem-solving schemas for specific tasks but also to abstract general meta-cognitive planning abilities, such as task decomposition, subgoal management, and dynamic strategy adjustment, enabling improved generalization or rapid adaptation to new tasks.
@@ 4.2.2. Prompt-based Methods
The model first formulates a complete action plan, subsequently followed by step-by-step execution by the system or the model itself. This mode offers clear logic but struggles to adapt to unexpected deviations during execution. For instance, the PAL  (Gao et al., 2023) method requires the model to write Python subroutines, which are then executed by an interpreter at runtime to obtain results. PoT  (Chen et al., 2023) further expresses the entire reasoning chain as program code, effectively decoupling the reasoning process from numerical computation. For multimodal and knowledge-intensive tasks, Chameleon  (Lu et al., 2023) explicitly separates planning and execution: an LLM acts as the planner, composing tool sequences (e.g., retrieval, Python functions, vision models, heuristic modules) as needed, which are then executed uniformly to produce the answer.
@@ 4.3.2. End-to-end Training
  • Trajectory-level credit assignment. Rewards are computed solely from the final, verifiable outcome and applied uniformly across the entire trajectory, implicitly treating all intermediate steps as equal contribution. Representative works include Search-R1 (Jin et al., 2025), R1-Searcher  (Song et al., 2025a), and ReSearch (Chen et al., 2025b), which optimize multi-turn retrieval with simple outcome-based rewards and use retrieval-token masking to limit overfit. AutoCoA (Zhang et al., 2025e) internalizes a Chain-of-Action agent: while its SFT stage contains step-level action triggers, its RL phase adopts trajectory-level credit assignment. ARTIST (Singh et al., 2025) unifies agentic reasoning, tool integration, and RL in a single framework where the model autonomously decides when, how, and which tools to invoke; optimization relies on outcome-based RL without step-level supervision.
@@
  • Step-level credit assignment. The goal is to identify which intermediate steps or turns substantially contribute to the final outcome. In multi-turn agent RL, finer-grained credit enables targeted improvement of specific actions rather than averaging rewards across all steps. For example, RAGEN (Wang et al., 2025f) shows that robust reasoning fails to emerge in multi-turn agent RL without fine-grained rewards, underscoring the need for granular attribution. Several lines of work pursue finer credit in multi-turn RL: Multi-Turn-RL (Zeng et al., 2025d) estimates advantages directly at the turn level; SPA-RL (Wang et al., 2025d) decomposes the terminal reward into per-step progress; GiGPO (Feng et al., 2025b) combines trajectory- and step-level signals without adding a critic or extra sampling, significantly improving multi-turn training; StepSearch (Wang et al., 2025i) frames retrieval and tool use through policy updates guided by information gain at each step, to mitigate policy drift from sparse global signals. Collectively, these methods shift reward from the terminal outcome to individual step, aligning gradients with effective actions and stabilizing training.
@@
  • Static/simulated environments.   These environments are fixed and typically preprocessed, yielding stable feedback and low noise. As a result, training is more stable than in real settings. For instance, ZeroSearch (Sun et al., 2025b) simulates a search engine with an LLM and applies curriculum RL to progressively degraded synthetic documents, substantially reducing external API costs and stochastic noise. MaskSearch (Wu et al., 2025c) develops “deep research” capabilities under controlled-noise conditions by using an offline retriever. ReSearch (Chen et al., 2025b) jointly optimizes reasoning and retrieval via RL and is commonly trained on offline or semi-online, controlled corpora. AutoCoA (Zhang et al., 2025e) internalizes a world model to reduce reliance on real-world interactions. In aggregate, this class of methods makes a pragmatic trade-off: sacrificing a degree of real-world adaptivity in exchange for the benefits of lower interaction costs, more stable training, and greater reproducibility.
@@
While a majority of studies integrate tools on non-reasoning base models (Jin et al., 2025; Wei et al., 2025c; Da et al., 2025), several recent end-to-end training studies (AutoCoA (Zhang et al., 2025e), WebDancer (Wu et al., 2025d), SkyRL (Cao et al., 2025), etc.) have begun to explore training on more capable LRMs. For instance, ReTool (Feng et al., 2025a) trains tool integration within the same tool-augmented RL framework using, respectively, a non-reasoning base (Qwen2.5-32B-Instruct (Qwen et al., 2025)) and a reasoning base (DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025)). The results show that, under matched training conditions, the reasoning base performs better, corroborating that base-model reasoning capacity enhances the learning of tool-integrated reasoning.

### v2

@@ 1.1. Paradigms
  • Memory has long been addressed with external modules. Short-term memory was typically maintained through conversation summary (Langchain Team, 2023), where long interaction histories were summarized and reinserted to fit within the context window. Long-term memory usually relied on retrieval-based methods, most prominently through Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), which stored past interactions in a vector database and retrieved them on demand.
@@
  • Memory has also begun to shift from external modules to model-native mechanisms. For short-term memory, Qwen-2.5-1M (Yang et al., 2025e) leveraged synthetic long-sequence data to extend the native context windows, addressing the challenge of “remembering” long-horizon information. Beyond remembering, recent works attempt to effectively “using” the information, e.g., MemAct (Zhang et al., 2025d) reframed context management as a tool the agent learns to call, proactively deciding when and how to store or retrieve information based on dynamic state and environmental feedback. For long-term memory, representative works like MemoryLLM (Wang et al., 2024a) pioneered parameterizing memory directly, where a set of latent memory tokens is continuously updated as part of the model’s forward pass, resulting in automatically updated internal knowledge.
@@ 1.2. Applications
The paradigm shifted with OpenAI’s introduction of the first model-native Deep Research agent (OpenAI, [n. d.]), which was fine-tuned based on its agentic foundation model o3. This approach, where the model’s internal policy learns to strategize the entire research process, significantly enhanced long-horizon consistency and the depth of information discovery, establishing Deep Research as an advanced assistant capable of tackling rigorous professional tasks. More recently, Tongyi Lab’s WebAgent series have advanced the open-source implementation of model-native Deep Research agents. This includes WebSailor (Li et al., 2025k), which addressed the critical challenge of synthesizing high-quality trajectory data for uncertain tasks, and culminated in the Tongyi DeepResearch model (Lab, 2025), which is capable of executing complex, multi-step research tasks in dynamic web environments.
@@ 1.3. Algorithms
To address these issues, a new set of outcome-driven RL algorithms has emerged to address the practical challenges of training stability and efficiency in long-horizon tasks. To counter the inefficiency of PPO’s large critic network, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) was proposed. It introduces a lightweight evaluation paradigm that computes advantages based on the relative rewards within a group of sampled responses, circumventing the need for an absolute value critic and improving training stability. Further refining this, Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO) (Yu et al., 2025b) improves performance in multi-turn interactions by decoupling the clipping mechanism for positive and negative advantages and employing dynamic sampling strategies, making it particularly effective for training long-horizon agents.
@@ 2.1. Necessity: The Shortage of Procedural Data
where $\mathcal{M}_{\text{base}}$ provides general world knowledge and reasoning priors, $\mathcal{A}_{\text{learn}}$ (e.g., RL or preference optimization) adapts and refines these capabilities through interaction and optimization, and $\mathcal{E}_{\text{task}}$ defines the environment, tool set, and reward signals that contextualize learning.
This formula effectively encapsulates the converging trend of recent AI research, particularly in agent training. Within this framework, contributions from the community tend to focus on one of several key facets, which are often deeply interrelated with the design of the “Task”:
  • Data Synthesis: Creating high-quality, large-scale interaction data to support the RL agent’s need for experience, enabling it to learn under diverse constraints.
  • Reward Function Design: Crafting sophisticated reward functions for complex tasks where outcomes are not easily verified, often balancing outcome-based rewards with process-based guidance.
  • Environment and Benchmark Construction: Building the simulation environments and reproducible benchmarks that provide the interactive, verifiable, and challenging scenarios necessary for both training and evaluation.
“Task” ($\mathcal{E}_{\text{task}}$) is not merely the final objective, but defines the entire learning world allowing the Learning Algorithm ($\mathcal{A}_{\text{learn}}$) to effectively optimize the Base Model ($\mathcal{M}_{\text{base}}$).
This trend toward a unified methodology is quite similar to the development of classical physics. Before Newton, physical subfields were fragmented: Celestial mechanics was governed by Kepler’s laws describing planetary motion, Terrestrial mechanics was pioneered by Galileo’s studies of falling bodies, and fields like Optics and Fluid dynamics had their own disparate empirical and geometric principles. Newton’s “Three Laws of Motion”, “Law of Universal Gravitation”, together with the powerful mathematical tool of Calculus, provided a universal framework that unified these experimental research under a single set of principles, achieving an unprecedented unification for Physics.
A similar dynamic can be observed in AI today with the emergence of the “LLM + RL” methodology. The LLM provides a unified model of world knowledge and foundational reasoning, analogous to a set of foundational principles. Meanwhile, RL offers a dynamic, goal-oriented optimization framework, analogous to a general-purpose problem-solving engine.
Following this parallel, the rise of this methodology “singularity” is shifting the focus of AI research, much as Newton shifted physics. In physics, this led to an expansion of applications into areas like fluid dynamics and celestial mechanics, a new focus on enhancing capabilities to solve complex challenges like the many-body problem, and eventually the theoretical innovations such as Quantum theory. Similarly, the practical focus for AI now ranges from the application expansion into practical domains like healthcare, scientific discovery, and social system simulation, to the capability enhancement required for challenges such as continual learning, safety, and alignment, and ultimately toward possibly new theories on general intelligence.
@@ 3.3.1. Supervised Learning
  • Tree search.  These typically combine Monte Carlo estimation or process reward models to enable fine-grained control over step-level quality (Li et al., 2025h, b). For instance, Marco-o1 (Zhao et al., 2024) employs Monte Carlo Tree Search (MCTS), using model confidence to guide exploration and synthesize high-quality reasoning data. WebSynthesis (Gao et al., 2025b) introduces a world model in Web-UI agent scenarios, combined with MCTS to generate web interaction trajectories at scale, supported by a two-stage curriculum learning strategy to enhance policy capability. ReST-MCTS* (Zhang et al., 2024g) integrates a process reward model with MCTS to collect higher-quality trajectories by assessing the value of each reasoning step. Regarding multimodal reasoning, AStar (Wu et al., 2025b) defines an executable visual action space and leverages MCTS to obtain high-quality visual reasoning data.
@@ 3.3.2. Reinforcement Learning
    To address scalability limitations of manual labeling, many studies employ Monte Carlo (MC) estimation or LLM-as-Judge (Li et al., 2025i) methods, sampling and comparing multiple reasoning paths to assess step quality. For example, Math-Shepherd (Wang et al., 2024b) assigns correctness scores to each mathimatical inference step via MC estimation, eliminating dependency on human annotation. OmegaPRM (Luo et al., 2024) enhances this approach by combining MC estimation with Monte Carlo Tree Search (MCTS), using binary search to locate the first erroneous step in reasoning chains, thereby improving data collection efficiency. Qwen-2.5-Math-PRM (Zhang et al., 2025g) notes that both MC estimation and LLM-as-Judge introduce noise in automatic labeling, while human annotation remains cost-prohibitive, thus proposing a hybrid approach to reduce noise and improve accuracy.
@@
Outcome reward, also referred to as outcome supervision, provides reward signals based solely on the correctness of the final answer. It offers a sparse yet unambiguous reward signal, grounded in the core hypothesis that a powerful foundation model can, through feedback on end results, autonomously explore and internalize effective reasoning policy. Exemplified by the RLVF approach adopted in DeepSeek R1 (DeepSeek-AI, 2025), outcome reward completely bypasses the complex evaluation of intermediate steps and only requires verifying the correctness of the final answer (Team, 2025; He et al., 2025a). In tasks with easily verifiable answers, such as multiple-choice questions and mathematical problems, RL is typically applied using verifiable outcome rewards to enhance the model’s reasoning capability (Trung et al., 2024; Zhang et al., 2024e; Luo et al., 2025c). Similarly, in the domain of code generation, methods like o1-Coder (Zhang et al., 2024b) and RLEF (Gehring et al., 2025) validate answer correctness based on program execution feedback.
@@ 3.4. Summary and Discussion
  • From supervised to self-supervised/unsupervised internalization: To reduce reliance on external reward signals, future research will explore self-supervised or unsupervised learning mechanisms. For example, by designing intrinsic reward signals such as exploration rewards based on information-theoretic principles like information gain or entropy (Zhang et al., 2025e), models could be motivated to autonomously discover novel and complex reasoning strategies, enabling continuous learning without explicit external objectives.
  • From single-task to general and transferable planning: Current models are often trained for planning within specific domains (e.g., mathematics or code). A significant future challenge lies in achieving cross-domain generalization and transfer of planning abilities (Su et al., 2025; Liu et al., 2025d). This requires models not only to learn problem-solving schemas for specific tasks but also to abstract general meta-cognitive planning abilities, such as task decomposition, subgoal management, and dynamic strategy adjustment, enabling improved generalization or rapid adaptation to new tasks.
@@ 4.2.2. Prompt-based Methods
The model first formulates a complete action plan, subsequently followed by step-by-step execution by the system or the model itself. This mode offers clear logic but struggles to adapt to unexpected deviations during execution. For instance, the PAL  (Gao et al., 2023) method requires the model to write Python subroutines, which are then executed by an interpreter at runtime to obtain results. PoT  (Chen et al., 2023) further expresses the entire reasoning chain as program code, effectively decoupling the reasoning process from numerical computation. For multimodal and knowledge-intensive tasks, Chameleon  (Lu et al., 2023b) explicitly separates planning and execution: an LLM acts as the planner, composing tool sequences (e.g., retrieval, Python functions, vision models, heuristic modules) as needed, which are then executed uniformly to produce the answer.
@@ 4.3.2. End-to-end Training
  • Trajectory-level credit assignment. Rewards are computed solely from the final, verifiable outcome and applied uniformly across the entire trajectory, implicitly treating all intermediate steps as equal contribution. Representative works include Search-R1 (Jin et al., 2025), R1-Searcher  (Song et al., 2025a), and ReSearch (Chen et al., 2025b), which optimize multi-turn retrieval with simple outcome-based rewards and use retrieval-token masking to limit overfit. AutoCoA (Zhang et al., 2025f) internalizes a Chain-of-Action agent: while its SFT stage contains step-level action triggers, its RL phase adopts trajectory-level credit assignment. ARTIST (Singh et al., 2025) unifies agentic reasoning, tool integration, and RL in a single framework where the model autonomously decides when, how, and which tools to invoke; optimization relies on outcome-based RL without step-level supervision.
@@
  • Step-level credit assignment. The goal is to identify which intermediate steps or turns substantially contribute to the final outcome. In multi-turn agent RL, finer-grained credit enables targeted improvement of specific actions rather than averaging rewards across all steps. For example, RAGEN (Wang et al., 2025g) shows that robust reasoning fails to emerge in multi-turn agent RL without fine-grained rewards, underscoring the need for granular attribution. Several lines of work pursue finer credit in multi-turn RL: Multi-Turn-RL (Zeng et al., 2025d) estimates advantages directly at the turn level; SPA-RL (Wang et al., 2025d) decomposes the terminal reward into per-step progress; GiGPO (Feng et al., 2025b) combines trajectory- and step-level signals without adding a critic or extra sampling, significantly improving multi-turn training; StepSearch (Wang et al., 2025j) frames retrieval and tool use through policy updates guided by information gain at each step, to mitigate policy drift from sparse global signals. Collectively, these methods shift reward from the terminal outcome to individual step, aligning gradients with effective actions and stabilizing training.
@@
  • Static/simulated environments.   These environments are fixed and typically preprocessed, yielding stable feedback and low noise. As a result, training is more stable than in real settings. For instance, ZeroSearch (Sun et al., 2025b) simulates a search engine with an LLM and applies curriculum RL to progressively degraded synthetic documents, substantially reducing external API costs and stochastic noise. MaskSearch (Wu et al., 2025c) develops “deep research” capabilities under controlled-noise conditions by using an offline retriever. ReSearch (Chen et al., 2025b) jointly optimizes reasoning and retrieval via RL and is commonly trained on offline or semi-online, controlled corpora. AutoCoA (Zhang et al., 2025f) internalizes a world model to reduce reliance on real-world interactions. In aggregate, this class of methods makes a pragmatic trade-off: sacrificing a degree of real-world adaptivity in exchange for the benefits of lower interaction costs, more stable training, and greater reproducibility.
@@
While a majority of studies integrate tools on non-reasoning base models (Jin et al., 2025; Wei et al., 2025c; Da et al., 2025), several recent end-to-end training studies (AutoCoA (Zhang et al., 2025f), WebDancer (Wu et al., 2025d), SkyRL (Cao et al., 2025), etc.) have begun to explore training on more capable LRMs. For instance, ReTool (Feng et al., 2025a) trains tool integration within the same tool-augmented RL framework using, respectively, a non-reasoning base (Qwen2.5-32B-Instruct (Yang et al., 2025e)) and a reasoning base (DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025)). The results show that, under matched training conditions, the reasoning base performs better, corroborating that base-model reasoning capacity enhances the learning of tool-integrated reasoning.